# Scalable Approximate Biclique Counting over Large Bipartite Graphs

### v1

@@ 5.1. Experimental Setting
Datasets. We use nine real datasets from different domains, which are available at SNAP (Project, 2009), Laboratory of Web Algorithmics (of Web Algorithmics, 2013), and Konect  (Konect, 2006). Table 1 shows the statistics of these graphs.
Baselines. We compare our method with the baselines from highly related works. We summarize the core ideas of each baseline method as follows.
@@
  • EP/Zz++ (Ye et al., 2023): the state-of-the-art algorithm for approximate biclique counting. It first partitions the graph into two regions: a dense region (a subgraph containing only high-degree vertices) and a sparse region (a subgraph containing only low-degree vertices). For the sparse region, EP/Zz++ utilizes EPivoter for exact counting, while for the dense region, it proposes a zigzag path-based sampling algorithm for approximate counting. Specifically, it leverages the fact that a $(p,q)$-biclique must contain a fixed number of $(min\{p,q\})$-zigzag paths. Based on this property, the algorithm first counts the number of $h$-zigzag paths and then samples $T$ such paths, where $h=min\{p,q\}$. Finally, it estimates the number of $h$-bicliques based on their proportion with the sampled paths.
  • CBS: our proposed approximation algorithm, which is introduced in Section 4.
Notice that we compare EP/Zz++ instead of EP/Zz (Ye et al., 2023), since the former one is a better version of the latter in terms of efficiency and accuracy. We implement all the algorithms in C++ and run experiments on a machine having an Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and 512GB of memory, with Ubuntu installed.
Parameter Settings. In our experiments, we following the existing work (Ye et al., 2023) setting $T$=$10^{5}$ as default values. We define the estimation error as $\frac{|\hat{C}-C|}{C}$, where $C$ denotes the exact count of bicliques, and $\hat{C}$ represents its approximated value. To ensure the reliability of our results, we run each approximation algorithm 10 times, with the reported error representing the average across these executions. For some datasets, the exact count of bicliques cannot be computed (e.g., ER and DE), we evaluate the estimation error differently by using the average approximated biclique count from 10 runs as the exact count (i.e., $C$).
@@ 5.2. Overall Comparison Results
1. Efficiency of All Algorithms. Figure 4 depicts the average running time of all the biclique counting algorithms on nine datasets for counting all $3\leq p,q\leq 9$ bicliques. We make the following observations and analysis: (1) Our method CBS is up to two orders of magnitude faster than all competitors, this is mainly because our algorithm has a better theoretical guarantee. (2) on almost all datasets, CBS is at least 10×10$\times$10 × faster than all methods, except DBLP dataset. On this dataset, BCList++ achieves the best performance, as graph reduction significantly reduces $|U|$ and $|V|$, allowing it to perform efficiently without extra initialization steps. Meanwhile, CBS still outperforms EPivoter and EP/Zz++. (3) On the two largest datasets, ER and DE, the exact algorithms fail to count the bicliques within $10^{5}$ seconds, while both approximate algorithms successfully complete the task. Our method, CBS, demonstrates at least 3 times faster performance compared to EP/Zz++.
@@
GHSOWutIMDBActor2ARDBLPERDE0010101010202020203030303040404040505050506060606070707070Dataseterror (%)EP/Zz++     CBS
@@
2. Accuracy of All Algorithms. Figure 5 illustrates the average error rates of CBS and EP/Zz++ across all datasets for biclique sizes ranging from 3 to 9 (i.e., $3\leq p,q\leq 9$). We can see that our algorithm demonstrates up to a 8×8$\times$8 × reduction in error compared to EP/Zz++, thanks to our carefully designed coloring scheme and $(p,q)$-broom-based sampling technique. Note that for the first three smaller datasets, EP/Zz++ exhibits significant errors, with at least 38.9% inaccuracy, rendering its results unreliable. In contrast, our method, CBS, maintains a maximum error rate of 16.9% under the same number of sampling rounds. Combining these observations with the performance results shown in Figure 4, we can conclude that our algorithm demonstrates an even greater advantage when considering the trade-off between accuracy and efficiency. This indicates that to achieve the same error rate, our algorithm would likely exhibit an even more substantial performance advantage over EP/Zz++.
@@
3. Effect of $p$ and $q$. Figures LABEL:figure:sampling_time_heatmap and LABEL:figure:error_heatmap illustrate the sampling time and error rate, respectively, of CBS and EP/Zz++ across various $p$ and $q$ values. In each figure, rows represent $p$ values and columns represent $q$ values. Each cell displays the sampling time (in Figure LABEL:figure:sampling_time_heatmap) or the estimation error (in Figure LABEL:figure:error_heatmap) for counting the corresponding $(p,q)$-bicliques. Clear, our method, CBS, outperforms EP/Zz++ by up to two orders of magnitude in both sampling time and accuracy. For instance, with $p=5$ and $q=3$ on the ER dataset, EP/Zz++ requires 822.6 seconds for sampling, whereas our algorithm completes this stage in just 5.3 seconds. In terms of accuracy, on the Wut dataset with $p=5$ and $q=7$, EP/Zz++ has an estimation error of 202.49, while our algorithm achieves a significantly lower error of 1.91. This observation aligns with our analysis in Section 4.
LABEL:Tcurve
LABEL:Errorcurve
4. Effect of $T$. We evaluate the effect of sample numbers on sampling time and error rate. The results on three datasets are shown in Figures LABEL:figure:Timecurve and LABEL:figure:T_curve. Based on these results, we observe the following: (1) For sampling time, our algorithm consistently outperforms EP/Zz++, particularly when the sample numbers are relatively low (e.g., from $10^{1}$ to $10^{3}$). (2) Our algorithm consistently produces more accurate results than EP/Zz++, regardless of whether the sample size is high or low. (3) Even with very few samples, our algorithm achieves acceptable solutions. For instance, on the AR dataset, CBS requires only $10^{3}$ samples to achieve a solution with a $30\%$ error rate, whereas EP/Zz++ requires over $10^{4}$ samples.
Detailed Analysis
In addition, more detailed analysis about CBS is provided in Appendix A, including Ablation Study, Time Cost of Different Stages, and Statistical of the Hyper-parameters. We only summarize the key conclusions here: (1) In high-accuracy scenarios, our coloring technique can significantly improve accuracy without highly impacting sampling time. (2) On more than half of the datasets, our algorithm requires less initializing and sampling time, compared to EP/Zz++. (3) When error rate $\epsilon$ is fixed, our algorithm CBS consistently requires fewer samples than EP/Zz++.
6. CONCLUSION
In this paper, we tackled the $(p,q)$-biclique counting problem in large-scale bipartite graphs, crucial for applications like recommendation systems and cohesive subgraph analysis. To address scalability and accuracy issues in existing methods, we proposed a novel sampling-based algorithm leveraging $(p,q)$-brooms, special spanning trees within $(p,q)$-bicliques. Utilizing graph coloring and dynamic programming, our method efficiently approximates $(p,q)$-biclique counts with unbiased estimates and provable error guarantees. Experimental results on nine real-world datasets show that our approach outperforms state-of-the-art methods, achieving up to 8$\times$ error reduction and 50$\times$ speed-up. Interesting future work includes extending our method to dynamic bipartite graphs with evolving structures and exploring its application to counting motifs/cliques in heterogeneous information networks.
Appendix A Detailed Analysis
@@
In this section, we extensively evaluate and analyze CBS from different angles.
A.1. Ablation Study
@@
To evaluate the effect of our coloring technique, we design a new variant of CBS by removing the vertex coloring step, denoted by BS. We then run them on two datasets and report the results in Figure 9. As we shall see, CBS generally achieves lower error with the same number of samples, aligning with our previous analysis. While the coloring technique slightly increases sampling time initially, the gap diminishes as sampling iterations increase. This demonstrates that in high-accuracy scenarios, our coloring technique can significantly improve accuracy without highly impacting sampling time.
@@
BS CBS BS CBS
$10^{2}$ 72.75 81.63 60.60 37.70
$10^{3}$ 43.33 28.47 22.03 23.50
$10^{4}$ 13.08 11.54 13.51 12.64
$10^{5}$ 4.37 3.35 5.11 4.59
$10^{6}$ 1.35 1.13 1.71 1.43
[…]
Figure 9. Results of BS and CBS with varying $T$.
A.2. Time Cost of Different Stages
For the two approximation algorithms, CBS and EP/Zz++, both require an initialization stage to precompute some auxiliary data for sampling. Specifically, in CBS, we need to assign a color number for each vertex (i.e., coloring) and count the number of $(p,q)$-brooms in bipartite graph, while in EP/Zz++, it needs to count the number of $h$-zigzag paths bipartite graph, where $h$ = $\min\{p,q\}$. In Figure  10, we report the initializing and sampling times for these algorithms across all datasets. We observe that on more than half of the datasets, our algorithm requires less initializing and sampling time, which indicates that the number of $(p,q)$-brooms in the bipartite graph is typically less than $h$-zigzag paths (using in EP/Zz++). In terms of sampling time, our algorithm is up to two orders of magnitude faster than EP/Zz++. While on two datasets, GH and ER, EP/Zz++ is 10 times faster in initializing, CBS achieves 100 times lower sampling time, making the total time (i.e., initializing plus sampling) of our algorithm still highly efficient.
@@ A.2. Time Cost of Different Stages
A.3. Statistical of the Hyper-parameters
Recall that in CBS and EP/Zz++, when $\epsilon$ is fixed, their required sample sizes are proportional to $\frac{B^{2}}{C^{2}}$ and $\frac{Z^{2}}{\rho^{2}}$, respectively. As shown in Table 2, we report the values of $\frac{B^{2}}{C^{2}}$ and $\frac{Z^{2}}{\rho^{2}}$ on the Amazon dataset for varying $p$ and $q$. Similar trends are observed across other datasets. Based on Table 2, our algorithm CBS consistently requires fewer samples than EP/Zz++, explaining why CBS achieves the same or even lower estimation error with a smaller sample size. For instance, CBS requires up to 88$\times$ fewer samples than EP/Zz++ on $p$ = 5 and $q$ = 9, demonstrating its efficiency in reducing sampling overhead while maintaining accuracy.

### v2

@@ 5.1. Experimental Setting
Datasets.
We use nine real datasets from different domains, which are available at SNAP (Project, 2009), Laboratory of Web Algorithmics (of Web Algorithmics, 2013), and Konect  (Konect, 2006). Table 1 shows the statistics of these graphs.
Methods.
We compare our method with the baselines from highly related works, summarized as follows:
@@ Methods.
  • EP/Zz++ (Ye et al., 2023): the state-of-the-art algorithm for approximate counting. It first partitions the graph into two regions: a dense region (a subgraph containing high-degree vertices) and a sparse region (a subgraph containing low-degree vertices). For the sparse region, EP/Zz++ utilizes EPivoter for exact counting, while for the dense region, it proposes a zigzag path-based sampling algorithm for approximate counting. Specifically, it leverages the fact that a $(p,q)$-biclique must contain a fixed number of $(min\{p,q\})$-zigzag paths. Based on this property, the algorithm first counts the number of $h$-zigzag paths and then samples $T$ such paths, where $h=min\{p,q\}$. Finally, it estimates the number of $h$-bicliques based on their proportion with the sampled paths.
  • CBS: our proposed approximation algorithm (Section 4).
Notice that we compare EP/Zz++ instead of EP/Zz (Ye et al., 2023), since the former one is a better version of the latter in terms of efficiency and accuracy. For completeness, we also conduct some initial experiment on EP/ZZ. Specifically, on the two largest datasets, ER and DE, EP/Zz failed to produce results within $10^{5}$ seconds for any values of $p$ and $q$ $(3\leq p,q\leq 9)$. Thus, it is not applicable to scalable datasets in practice. Thus, we focus our evaluation primarily on EP/Zz++ in our main experiments.
There is also another potential baseline introduced in (Ye et al., 2024). The method takes over 1000 seconds to run on the DBLP dataset, while our algorithm requires only around 1 second. We have performed a primary initial evaluation based on the conclusions reported in the article. They claim to outperform BCList++ and EPivoter by two orders of magnitude in runtime. However, both baselines fail to return results on the ER and DE datasets within $10^{5}$ seconds. Even based on their own estimation, their method remains significantly slower than ours in practice. As a result, we do not include it as a baseline in our empirical study. We implement all the algorithms in C++ and run experiments on a machine having an Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and 512GB of memory, with Ubuntu installed.
Parameter Settings.
[…]
@@ 5.2. Overall Comparison Results
1. Efficiency of All Algorithms.
Figure 4 depicts the average running time of all the biclique counting algorithms on nine datasets for counting all $3\leq p,q\leq 9$ bicliques. We make the following observations and analysis:
  1. (1)
    Our method CBS is up to two orders of magnitude faster than all competitors, this is mainly because our algorithm has a better theoretical guarantee.
  2. (2)
    On almost all datasets, CBS is at least $10\times$ faster than all methods, except DBLP dataset. On this dataset, BCList++ achieves the best performance, as graph reduction significantly reduces $|U|$ and $|V|$, allowing it to perform efficiently without extra steps. Meanwhile, CBS still outperforms EPivoter and EP/Zz++.
  3. (3)
    On the two largest datasets, ER and DE, the exact algorithms fail to terminate in $10^{5}$ seconds, while both approximate algorithms successfully complete the task. Our method, CBS, demonstrates at least 3 times faster performance compared to EP/Zz++.
Note that besides Figure 4, we choose to simply compare the sampling time with EP/Zz++ in Figure 6 and Figure 8, which is for exposing the significant improvement in the sampling efficiency.
@@
GHSOWutIMDBActor2ARDBLPERDE01010202030304040505060607070Dataseterror (%)EP/Zz++  CBS
@@
Figure 5 illustrates the average error rates of CBS and EP/Zz++ across all datasets for biclique sizes ranging from 3 to 9 (i.e., $3\leq p,q\leq 9$). We can see that our algorithm demonstrates up to a $8\times$ reduction in error compared to EP/Zz++, thanks to our carefully designed coloring scheme and $(p,q)$-broom-based sampling technique. Note that for the first three smaller datasets, EP/Zz++ exhibits significant errors, with at least 38.9% inaccuracy, rendering its results unreliable. In contrast, our method, CBS, maintains a maximum error rate of 16.9% under the same number of sampling rounds. Our algorithm also consistently obtains an outstanding accuracy in larger datasets. A minor exception occurs in ER that our algorithm obtains a slightly worse error ratio in only one case $(p,q)=(4,9)$. To compensate for this, since our sampling time is more than 70 times faster compared to EP/Zz++ (Figure 6 (a)(b)), by simply increasing the sampling times, we can decrease the error ratio while still outperforming the baseline. Initially, we sampled $10^{5}$ times. By sampling $5\times 10^{5}$, $10^{6}$ and $7\times 10^{6}$, our error ratio decreases to $1.06\%$, $1.01\%$, $0.34\%$, respectively. These results all outperform EP/Zz++ in both error ratio and sampling time.
Combining these observations with the performance results shown in Figure 4, we can conclude that our algorithm demonstrates an even greater advantage when considering the trade-off between accuracy and efficiency. This indicates that to achieve the same error rate, our algorithm would likely exhibit an even more substantial performance advantage over EP/Zz++.
(a) ER, EP/Zz++
(b) ER, CBS
(c) DE, EP/Zz++
(d) DE, CBS
Figure 6. The heat-map of sampling time of EP/Zz++ and CBS with varying $p$ and $q$ (s).
(a) Wut, EP/Zz++
(b) Wut, CBS
(c) AR, EP/Zz++
(d) AR, CBS
Figure 7. The heat-map of estimation errors of EP/Zz++ and CBS with varying $p$ and $q$ (%).
3. Effect of $p$ and $q$.
@@
Figures 6 and 7 illustrate the sampling time and error rate, respectively, of CBS and EP/Zz++ across various $p$ and $q$ values. In each figure, rows represent $p$ values and columns represent $q$ values. Each cell displays the sampling time (in Figure 6) or the estimation error (in Figure 7) for counting the corresponding $(p,q)$-bicliques. Clearly, our method, CBS, outperforms EP/Zz++ by up to two orders of magnitude in both sampling time and accuracy. For instance, with $p=5$ and $q=3$ on the ER dataset, EP/Zz++ requires 822.6 seconds for sampling, whereas our algorithm completes this stage in just 5.3 seconds.
In terms of accuracy, on the Wut dataset with $p=5$ and $q=7$, EP/Zz++ has an estimation error of 202.49, while our algorithm achieves a significantly lower error of 1.91. This observation aligns with our analysis in Section 4. As a minor observation in Figure 6, the sampling time of CBS is more affected by q than by p while EP/Zz++ does not exhibit this characteristic. This observation indeed demonstrates the difference between Ep/Zz++ and our algorithm to some levels. Specifically, EP/Zz++ needs to sample a zig-zag path with $h=\min(p,q)$ to approximate the number of $(p,q)$-bicliques, therefore its sampling time increases in a staircase manner w.r.t $\min(p,q)$. On the other hand, our algorithm samples brooms, which is different from EP/Zz++, this could be why we exhibit different characteristics. In our case, one of the possible reasons is that we start ”growing” a broom from the left side during the sampling process.
$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{6}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$Tsampling time(s)
(a) ER
$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{6}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{3}$Tsampling time(s)
(b) AR
Figure 8. Average sampling time of EP/Zz++ and CBS with varying $T$.
$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$$0^{6}11$3310103030100100Terror(%)
(a) Wut
$10^{2}$$10^{3}$$10^{4}$$10^{5}$$0^{6}11$3310103030100100Terror(%)
(b) AR
Figure 9. Average error of EP/Zz++ and CBS with varying $T$.
4. Effect of $T$.
We evaluate the effect of sample numbers on sampling time and error rate. The results on three datasets are shown in Figures 8 and 9. Based on these results, we observe the following:
  1. (1)
    For sampling time, our algorithm consistently outperforms EP/Zz++, particularly when the sample numbers are relatively low (e.g., from $10^{1}$ to $10^{3}$).
  2. (2)
    Our algorithm consistently produces more accurate results than EP/Zz++, regardless of whether the sample size is high or low.
  3. (3)
    Even with very few samples, our algorithm achieves acceptable solutions. For instance, on the AR dataset, CBS requires only $10^{3}$ samples to achieve a solution with a $30\%$ error rate, whereas EP/Zz++ requires over $10^{4}$ samples.
6. Detailed Analysis
@@
In this section, to provide a better understanding of our algorithm, besides the primary evaluation in Section 5, we now extensively evaluate and analyze CBS from different angles:
$\bullet$ An ablation study on the effect of our coloring technique, showing that it can significantly improve accuracy without highly impacting sampling time.
$\bullet$ An ablation study on the coloring technique conducted using different ordering strategies.
$\bullet$ A study on the time cost of each step in CBS, including the initialization and sampling time, compared to the baseline EP/Zz++.
$\bullet$ A study on the statistical of the hyper-parameters that determines the error ratio, which provides evidence that our algorithm CBS consistently requires fewer samples than EP/Zz++.
$\bullet$ A study of the memory consumption of our algorithm.
$\bullet$ An additional scalability experiment conducted by randomly sampling different percentages of vertices in the same graph.
$\bullet$ An additional efficiency experiment with small $p,q$ aligning with the empirical study in (Ye et al., 2023).
6.1. Ablation Study on the Coloring Technique
We conduct an ablation study on the effect of our coloring technique. To begin with, in order to evaluate the effect of our coloring technique, we design a new variant of CBS by removing the vertex coloring step, denoted by BS. Then, note that in our proposed coloring methods, we decide to traverse the neighbors in a random order manner. A question is raised that if we follow different order of neighbors, would the performance be significantly affect? To answer this question, we design three more variants of CBS by using two different ordering strategies in coloring: the ascending order and descending order of degrees, denoting as ACBS and DCBS.
@@ 6.1. Ablation Study on the Coloring Technique
To compare the inaccuracy, we first run these variants on the two datasets AR and Wut with varying parameters of $T$, and report the results in Figure 10. Compared to BS (without the coloring), CBS generally achieves lower error with the same number of samples, aligning with our previous analysis. Although the coloring technique initially slightly increases the sampling time, the gap diminishes as the number of sampling iterations increases. This demonstrates that in high-accuracy scenarios, our coloring technique can significantly enhance accuracy without significantly impacting the sampling time. Compared to ACBS and DCBS, there is no significant difference in error in the three strategies. This confirms that our algorithm is robust to the choice of coloring order.
In Figure 11, we run all four algorithms on the dataset ER with varying values of $T$ to compare their efficiency. We can see that without the coloring (BS), the total runtime does not decrease much. On the other hand, the average runtime of CBS, ACBS, and DCBS is similar. This again confirms the robustness of our coloring method. Combining with the accuracy result, we can see that the coloring method is effective, as with a slight increase in runtime, the accuracy is enhanced. Overall, our ablation study demonstrates that the coloring process is vital in counting bicliques.
@@
BS CBS ACBS DCBS BS CBS ACBS DCBS
$10^{2}$ 72.75 81.63 98.50 78.99 60.60 37.70 87.07 33.91
$10^{3}$ 43.33 28.47 32.44 32.43 22.03 23.50 24.48 21.16
$10^{4}$ 13.08 11.54 11.55 10.62 13.51 12.64 13.41 12.89
$10^{5}$ 4.37 3.35 3.92 3.40 5.11 4.59 4.44 3.98
$10^{6}$ 1.35 1.13 1.19 1.06 1.71 1.43 1.67 1.42
Figure 10. Average error of different coloring techniques in AR and Wut with varying $T$.
11
$10^{2}$$10^{3}$$10^{4}$$10^{5}$$10^{6}$5050100100200200400400800800TTtotal time (s)
Figure 11. Average runtime of different coloring techniques in ER with varying $T$.
6.2. Time Cost of Different Stages
The two approximation algorithms, CBS and EP/Zz++, both require an initialization stage to precompute some auxiliary data for sampling. Specifically, in CBS, we need to assign a color number for each vertex (i.e., coloring) and count the number of $(p,q)$-brooms in bipartite graph, while in EP/Zz++, it needs to count the number of $h$-zigzag paths bipartite graph, where $h$ = $\min\{p,q\}$. In Figure  12, we report the initializing and sampling times for these algorithms across all datasets. We observe that on more than half of the datasets, our algorithm requires less initializing and sampling time, which indicates that the number of $(p,q)$-brooms in the bipartite graph is typically less than $h$-zigzag paths (using in EP/Zz++). In terms of sampling time, our algorithm is up to two orders of magnitude faster than EP/Zz++. While on two datasets, GH and ER, EP/Zz++ is 10 times faster in initializing, CBS achieves 100 times lower sampling time, making the total time (i.e., initializing plus sampling) of our algorithm still highly efficient.
@@ 6.2. Time Cost of Different Stages
6.3. Statistical of the Hyper-parameters
[…]
@@ 6.3. Statistical of the Hyper-parameters
6.4. Memory Consumption
The theoretical guarantee of the memory usage is given in Theorem 4.8. We now present the actual memory consumption on each dataset in Table 3 (in Megabytes). Specifically, we report the average memory usage on each dataset for $3\leq p,q\leq 9$. As we can observe, the memory usage is similar to all other baselines for most datasets. However, our memory consumption increases for the two large datasets. This is in line with our analysis, as it is dominated by $(p+q)|E|$. Our algorithm requires to store the intermediate result for retrieving the broom structure. In result, the memory is mainly for storing the dynamic programming table.
Table 3. Memory consumption of different counting algorithms (in MBs).
dataset BCList++ EPivoter EP/Zz++ CBS
GH 21.3 880.8 21.3 462.1
SO 40.9 901.0 73.7 120.8
Wut 52.2 911.0 69.8 312.8
IMDB 66.5 926.0 106.1 164.2
Actor2 88.3 946.9 103.1 268.5
AR 152.7 1004.5 295.8 461.0
DBLP 319.5 1166.3 444.5 718.2
ER X X 402.7 46600.6
DE X X 1744.8 195153.0
6.5. Scalability Experiment
We consider counting bicliques approximately, mainly because all exact counting methods are highly inefficient in large graph data. Therefore, it is essential to verify the scalability of our algorithm thoroughly. Besides the main experiment in Section 5, we now conduct an additional graph scalability experiment that evaluates performance under dynamic modifications, which is done by randomly removing a percentage of edges. Specifically, we construct several graphs from DE as follows:
  1. (1)
    Randomly sample $60\%$, $80\%$, $100\%$ of the vertices on both sides and keep the induced graph.
  2. (2)
    Execute our algorithm on $p=6,q=8,T=10^{5}$.
  3. (3)
    Repeat (1) and (2) 10 times, compute the average runtime for initialization and sampling.
The three constructed graphs are DE60, DE80, and DE100. Shown in Table 4, we can observe that as the size of the graph increases, the sampling time also increases linearly while the error ratio is within a certain range. This shows that CBS’s scalability, which is expected to be particularly effective on large datasets.
Table 4. Scalability experiment.
init time(s) sample time(s) error(%)
DE60% 298.6 40.8 3.76
DE80% 1216.7 95.3 2.49
DE100% 2478.1 135.5 3.19
6.6. Efficiency Comparison on Small $p,q$
One of our major baselines, EPivoter is designed for exact counting on small $p,q$ (e.g., $p,q\leq 5$). However, our major comparison experiment is conducted for all combinations of $p,q\leq 9$. To bridge the gap, we provide an additional experiment using the same parameter settings as EPivoter, i.e., $p,q\leq 5$. Shown in Table 5, for ER and DE, EPivoter can not terminate in $10^{5}$ seconds. For other datasets, we observe that our method maintains its performance advantage even when $p$ and $q$ are small, highlighting its robustness to parameter settings. These additional results confirm that our method performs consistently well regardless of the parameter choices.
Table 5. Average runtime (s) of EPivoter and CBS for all $3\leq p,q\leq 5$.
Dataset EPivoter CBS
GH 805.4 3.4
SO 95.0 1.4
Wut 2429.4 4.3
IMDB 18.4 1.6
Actor2 53.4 2.8
AR 180.6 3.7
DBLP 18.5 1.9
ER INF 364.4
DE INF 3782.1
7. CONCLUSION
[…]
Acknowledgements.
Chenhao Ma was partially supported by NSFC under Grant 62302421, Basic and Applied Basic Research Fund in Guangdong Province under Grant 2025A1515010439, and the Guangdong Provincial Key Laboratory of Big Data Computing, The Chinese University of Hong Kong, Shenzhen. Weinuo Li and Can Wang were partially supported by the National Natural Science Foundation of China (No. 62476244), the Starry Night Science Fund of Zhejiang University Shanghai Institute for Advanced Study, China (Grant No: SN-ZJU-SIAS-001). Parts of this work was done while Jingbang was still at University of Waterloo.